a stone is thrown upwards from the ground with a velocity of 40m/s. how long does it take to reach its highest point
1
Expert's answer
2018-04-29T13:47:07-0400
At highest point velocity of the stone is zero
v_f=v_i-gt=0
So
t=v_i/g=40/10=4 s